OpenStudy (anonymous):

order of operations 3/4 divided by (1/3+1/2)

OpenStudy (rina.r):

order of operations is BIMDAS so i will solve the bracket first (1/3+1/2) 0.3 + 0.5 then 3/4 is 0.75 =0.8+0.75 =1.55 i think that's how it is.

OpenStudy (anonymous):

Unless they use PEMDAS, which in my opinion is a lot easier than BIMDAS, for the simple fact that I had to Google the word indices, and apparently it is the plural form of index. Now an index is a guide. Parentheses (Brackets) Exponents (indices) Multiplication Division Addition Subtraction Either one works, I guess it depends on where they live. I thought I'd throw both of them out there in case they use PEMDAS and have no idea what BIMDAS is. O.o